OpenStudy (anonymous):

Where does the x^2y’ term come from in 2xy + x^2y’ + 3y^2y’ + 2x = 0? By the way, 2xy + x^2y’ + 3y^2y’ + 2x = 0 is achieved after using implicit differentiation on x^2y + y^3 + x^2 = 8.

OpenStudy (anonymous):

\[x^2y+y^3+x^2=8\] take differentiate both sides the first term is = \(2xy + x^2 y'\) (this is a product, right?) second term is = \(3y^2y'\) the third one is = 2x --------------------------------------------------- \(2xy+x^2y'+3y^2y'+2x=0\)

OpenStudy (anonymous):

does it make sense??

OpenStudy (anonymous):

\[(x^2\color{red}{y})'= (x^2)'\color{red}{y}+x^2\color{red}{(y)'}=2xy+x^2y' \] Is it not clear enough??

OpenStudy (abb0t):

@charlotte123

OpenStudy (charlotte123):

@abb0t ;-;

OpenStudy (charlotte123):

T_T